OpenStudy (anonymous):

f(x) = (x-2)/(x+3) show that x is continuous at x=1

OpenStudy (anonymous):

Do I need to show that lim x->1 (f(x)) == f(1) ?

OpenStudy (anonymous):

yes. lim[x->1] f(x)=f(1) and that lim[x->1] f(x) exists, saying that both sides are equal. and 3rd condition, that f(1) exists, but once you get the first 1, the last 2 are obvious.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{x-2}{x+3}=f(1)\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{(x)-2}{(x)+3}=\frac{(1)-2}{(1)+3}=\frac{-1}{4}=-1/4\]

OpenStudy (anonymous):

so, you can ell right away that lim{x->1} f(x) is equal to f(1), since there are no restrictions for x to equal 1. it is continuous, fo all numbers besides x=-3, right? (vertical asymptote.)

OpenStudy (anonymous):

Sounds right, the prof wants us to show by comparing the limit to f(x) I think. How would you prove if there were a jump in the graph? Would you have to show that it is different from the left and right, as well as at the point itself?